OpenStudy (anonymous):

Really hard trig question?? Fan and medal please! (1-cosΘ)(1+cosΘ)/(1-sinΘ)(1+sinΘ)?

OpenStudy (anonymous):

Possible Answers: sin2 Θ cos2 Θ tan2 Θ cosine theta over sine theta

Nnesha (nnesha):

foil o^_^o denominator and numerator

OpenStudy (anonymous):

How do I do that?

OpenStudy (thomas5267):

\[ \frac{(1-\cos(\theta))(1+\cos(\theta))}{(1-\sin(\theta))(1+\sin(\theta))}=\frac{1-\cos^2(\theta)}{1-\sin^2(\theta)} \]

Nnesha (nnesha):

 bec one parentheses is negative and one is positive so square of first term and square of 2nd term middle terms would cancel out